### v1

# C-3DPO: Constrained Controlled Classification for Direct Preference Optimization

Abstract
Direct preference optimization (DPO)-style algorithms have emerged as a promising approach for solving the alignment problem in AI. We present a novel perspective that formulates these algorithms as implicit classification algorithms. This classification framework enables us to recover many variants of DPO-style algorithms by choosing appropriate classification labels and loss functions. We then leverage this classification framework to demonstrate that the underlying problem solved in these algorithms is under-specified, making them susceptible to probability collapse of the winner-loser responses. We address this by proposing a set of constraints designed to control the movement of probability mass between the winner and loser in the reference and target policies. Our resulting algorithm, which we call Constrained Controlled Classification DPO (C-3DPO), has a meaningful RLHF interpretation. By hedging against probability collapse, C-3DPO provides practical improvements over vanilla DPO when aligning several large language models using standard preference datasets.
Machine Learning, ICML
1 Introduction
The problem of ensuring that AI systems act in accordance with human preferences, also known as the alignment problem, has become a critical focus in machine learning. Reinforcement Learning from Human Feedback (RLHF) has emerged as a promising approach towards solving this problem (Christiano et al., 2017). RLHF proceeds by first learning a reward model (RM) from a winner-loser preference dataset followed by employing standard RL algorithms to maximize the learned RM while keeping the model close to a reference model.
Recent years have witnessed the emergence of algorithms that solve the two RLHF sub-problems as a single optimization problem. Chief among these methods is the Direct Preference Optimization (DPO) algorithm (Rafailov et al., 2023). It proceeds by leveraging the closed-form solution of the RLHF objective and using the preference dataset to align the model, thus obviating the explicit reward-learning phase of RLHF. Since then, numerous extensions and successors of DPO have been proposed, highlighting the need for a deeper examination of this new class of algorithms in order to unify and connect their underlying principles, and to highlight and fix their potential pitfalls.
In this paper, we first present a classification framework that reformulates the family of DPO-style algorithms. The primary benefit of this framework is that it unifies and extends a variety of existing DPO-style algorithms. Most notably, with specific choices of classification labels and loss function, we recover the popular DPO (Rafailov et al., 2023) and IPO (Azar et al., 2024) algorithms, revealing them as special cases of our classification formulation. Moreover, unlike traditional RLHF approaches that rely solely on binary preference pairs, our classification framework can naturally incorporate richer information, such as ranked lists of preferences and auxiliary score information (e.g., ratings) about the responses.
We next leverage this classification perspective to investigate a peculiar behavior observed in several DPO-style algorithms: the decrease, or even collapse to zero, of the winner-loser probabilities. Note that prior work has documented this phenomenon (Adler et al., 2024; Xu et al., 2024; Pal et al., 2024; Fisch et al., 2024; Xiao et al., 2024; Shen et al., 2024; Wu et al., 2024; D’Oosterlinck et al., 2024). Using our proposed classification framework, we pinpoint the root cause of this behavior: minimizing the loss in the DPO-style of algorithms only provides a single constraint for learning two probabilities, leaving the problem under-constrained. Leaning on this insight, we then propose a new set of constraints that can provably fix the under-specified nature of the DPO-style algorithms, thus effectively addressing the probability-collapse issue. These constraints are designed to control the movement of winner-loser probability mass between the reference and target policies. Our proposed algorithm, Constrained Controlled Classification DPO (C-3DPO), optimizes the DPO objective under these constraints.
We evaluate the effectiveness of our set of constraint in enhancing preference alignment. Across two standard datasets and three with up to 13B parameters, C-3DPO outperforms vanilla DPO and several other baselines, delivering higher-quality final models when assessed holistically on the standard MT Benchmark (Zheng et al., 2023).
2 Preliminaries
In this section, we present the key ingredients of preference optimization, which we will build upon in the subsequent sections. In this setting, we are given a dataset $\mathcal{D}$ of triplets $(x,y_{w},y_{l})$, where $x$ is a prompt, while $y_{w}$ and $y_{l}$ reflect our preference in choosing $y_{w}$ over $y_{l}$ conditioned on $x$. We are also given a reference $\pi_{\textrm{ref}}$ serving as a guardrail.
In RLHF, we first employ $\mathcal{D}$ to train a parameterized RM, $r_{\phi}$, and then use it to solve the following:
$\!\!\!\!\max_{\theta}\ \mathbb{E}_{x}\Big{[}\mathbb{E}_{y\sim\pi_{\theta}}\big{[}r_{\phi}(x,y)\big{]}-\beta\mathbb{KL}\big{(}\pi_{\theta}(\cdot|x)||\pi_{\textrm{ref}}(\cdot|x)\big{)}\Big{]},$ (1)
where $\beta>0$ is a hyper-parameter denoting the relative importance of reward maximization against ensuring a low deviation from $\pi_{\textrm{ref}}$. The RM is typically learned assuming that preferences follow the Bradley-Terry (BT) model:
$p(y_{w}\succ y_{l}\mid x)=\sigma\big{(}r(x,y_{w})-r(x,y_{l})\big{)}\;,$ (2)
where $\sigma(x)=1/(1+\exp(-x))$ is the sigmoid function and $r$ is the latent reward of the annotator. Therefore, fine-tuning $\pi_{\theta}$ in the RLHF approach is split into two stages: reward-learning using the BT model, followed by a policy optimization using (1).
More recently, a family of algorithms have emerged that solve the above two problems in a single stage, which is referred to as Direct Preference Optimization (DPO)-style algorithms. The key insight in DPO-style algorithms is that problem (1) admits the following closed-form solution:
$\pi^{*}(y|x)=\frac{\pi_{\textrm{ref}}(y|x)\exp\big{(}r(x,y)/\beta\big{)}}{Z(x)}\ ,$ (3)
where $Z(x)$ is the partition function and is generally intractable. We can rewrite (3) as
$r(x,y)=\beta\log\frac{Z(x)\pi^{*}(y|x)}{\pi_{\textrm{ref}}(y|x)}\ .$ (4)
Substituting $r(x,y)$ from (4), $Z(x)$ cancels out, leading to the optimization problem solved by DPO:
$\min_{\theta}\sum_{(x,y_{w},y_{l})\in\mathcal{D}}-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\;.$ (5)
3 A Classification View of DPO-style Algorithms
In this section, we show that DPO-style algorithms can be viewed as classification algorithms. The standard classification setting has three main ingredients.
First, we typically construct a classifier hypothesis space by defining a parametric form for probabilities assigned to each class, and then train the classifier to adjust these probabilities. We show that in DPO-style algorithms these probabilities are implicitly defined as follows:
$p_{\theta}(x,y_{w},y_{l})\mathrel{\mathop{:}}=\textrm{softmax}\big{(}r_{\theta}(x,y_{w}),r_{\theta}(x,y_{l})\big{)}\ ,$ (6)
where $r_{\theta}$ is the reward defined in (4) having substituted $\pi^{*}$ with $\pi_{\theta}$. It is straightforward to see that under distribution $p_{\theta}$, the probability assigned to the winner (preferred) response $y_{w}$, denoted by $p^{w}_{\theta}$, does not depend on the partition function $Z(x)$ and can be written as
$p^{w}_{\theta}(x,y_{w},y_{l}):=\frac{\big{(}\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\textrm{ref}}(y_{w}\mid x)}\big{)}^{\beta}}{\big{(}\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\textrm{ref}}(y_{w}\mid x)}\big{)}^{\beta}+\big{(}\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\textrm{ref}}(y_{l}\mid x)}\big{)}^{\beta}}\ .$ (7)
The probability assigned to the loser (dispreferred) response $y_{l}$, denoted by $p^{l}_{\theta}$, is defined similarly. Note that the distribution $p_{\theta}$ in (6) is a function of $\pi_{\theta}$, $\beta$, and $\pi_{\textrm{ref}}$, and can be thought of as a generalization of the conditional probability of a response $y$ given that $y\in\{y_{w},y_{l}\}$.
Second, in the standard classification setting, the dataset gives us access to labels, which we use to extract the probabilities associated with each class. To obtain these target probabilities, we simply use any distribution $p=(p^{w},p^{l})$ from the simplex $\Delta_{2}$, which is defined as the set of all vectors $p$ in $\mathbb{R}^{2}$ satisfying $p^{w},p^{l}\geq 0$ and $p^{w}+p^{l}=1$. In the most basic case we just use the one-hot vector $(p^{w},p^{l})=(+1,0)$ akin to using hard labels. This indicates that we desire to always prefer the answer $y_{w}$ to $y_{l}$ when presented with a choice between the two. More generally, we can use soft labels, meaning that we put some non-zero weight behind each class. Soft labels have shown to be quite useful in classification (Müller et al., 2019), and by extension have been useful in preference optimization where we are less certain about the given preference $y_{w}\succ y_{l}$.
Third, to solve a classification problem, we often define a differentiable loss function measuring the discrepancy between a distribution $p_{\theta}$ in the hypothesis space and the target distribution $p$. We can use any standard loss $\mathcal{L}$ for classification leading to the optimization problem: $\min_{\theta}\sum_{\mathcal{D}}{\mathcal{L}}(p_{\theta},p)$. The loss function should be non-negative $\mathcal{L}(p_{1},p_{2})\geq 0$, and satisfy $\mathcal{L}(p_{1},p_{2})=0$ if and only if $p_{1}=p_{2}$. A good example is the CE loss.
We now show that a large number of DPO-style algorithms can be viewed as specific instances of this classification framework. The generality arises from the ability to use (a) hard or soft labels for the target distribution $p$ and (b) different classification losses $\mathcal{L}$. To start we show that the popular DPO algorithm (Rafailov et al., 2023) can be recovered using hard labeling and the CE loss.
Remark 3.1 (DPO).
Suppose that we use the CE loss and hard labels $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(+1,0)$ in the classification framework above. Then, using (7), we can write
${\mathcal{L}}\big{(}p_{\theta},\ p\big{)}$ $=-\left(p^{w}\log p^{w}_{\theta}+p^{l}\log p^{l}_{\theta}\right)\;\;=\;-\log p^{w}_{\theta}$
$=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\ ,$
which is exactly the DPO loss (5) summed over $\mathcal{D}$.
Another popular DPO-style algorithm is IPO (Azar et al., 2024). While the derivation of IPO in the original paper looks completely different than DPO, we now show that IPO can also be viewed as a specific instance of our classification framework. In contrast to DPO, to obtain IPO we use soft labeling in conjunction with a less common loss function which we present below.
Remark 3.2 (IPO).
Suppose that we use the following loss
$\mathcal{L}\big{(}p_{\theta},\ p\big{)}=\big{(}\log(p^{w}_{\theta}/p^{l}_{\theta})-\log(p^{w}/p^{l})\big{)}^{2}\ ,$ (8)
and soft labels $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(\sigma(1/2),\sigma(-1/2))$ in the above classification framework. We will then recover the IPO loss (Eq. 17 in Azar et al. 2024). See Appendix A for more details.
Azar et al. (2024) motivated IPO primarily as a mean to overcome the issue of vanishing KL penalty in the DPO algorithm. We highlight that IPO can hedge against this issue merely because it employs soft classification labels. Stated differently, to avoid the vanishing KL, we do not have to commit to the specific IPO loss. We can achieve this desirable property using soft classification labels and for instance, the CE loss, instead of the square loss in IPO. In this case, we obtain the Conservative DPO (CDPO) algorithm (Mitchell, 2024) in which $p:=(p^{w},p^{l})=(1-\varepsilon,\varepsilon)$ for $20<\varepsilon<1/$ 2 (see more details in Appendix B). Similarly, to IPO, CDPO is free from the vanishing KL issue.
In Table 1, we show how several DPO-style algorithms can be formulated using our classification framework.
Data Algorithm Labels Loss
Pairs DPO(BT) Hard CE
CDPO Soft CE
IPO Soft Eq. (8)
List (Sec. 3.1) DPO(PL) Hard CE
Auxiliary Info RPO Soft CE
(Sec. 3.2) Distilled DPO Soft Eq. (8)
Table 1: Examples of DPO-style algorithms as classification. By DPO(BT) and DPO(PL), we refer to the DPO algorithm with Bradley-Terry and Plackett-Luce (see Sec. 3.1) models, respectively. More details on RPO (Adler et al., 2024) and Distilled DPO (Fisch et al., 2024) algorithms can be found in Sec. 3.2.
We now provide two straightforward extensions of the above classification framework that allow us to incorporate a larger set of existing DPO-style algorithms.
3.1 List of Preferences
Our classification framework can be extended to work with lists, rather than pairs, of preferences. In particular, assume that we have $N$ responses for each prompt $x$, giving us a dataset of the form $\mathcal{D}=\{(x,y_{1},y_{2},\ldots,y_{N})\}.$ In this case, we can define a list version of the probability vector $p_{\theta}(x,y_{1},\ldots,y_{N})$ similarly to (6), together with a target distribution $p$. With this simple extension we can now incorporate existing DPO-style algorithms that work with lists. For instance, we can show that DPO with the Plackett-Luce model for preferences (Rafailov et al., 2023) can be captured in our classification framework, again using hard labels and the CE loss. See Appendix C for a proof.
3.2 Auxiliary Information
A second important extension pertains to the definition of soft labels in our classification setting. So far we have only worked with soft labels that are fixed across the entire dataset, for instance, $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(\sigma(1/2),\sigma(-1/2))$ for all $(x,y_{w},y_{l})$ in IPO. These fixed labels are agnostic about any extra information we may have about our data triplets $(x,y_{w},y_{l})$. However, in some applications we may have access to some auxiliary scores, $s_{w},s_{l}$, (e.g., ratings) associated with each response, which can then be used to enrich our soft labels.
More formally, suppose now that our dataset is comprised of ${\mathcal{D}}=(x,y_{w},s_{w},y_{l},s_{l})$. To obtain the soft labels we can employ, for instance, $p=\text{softmax}(s_{w},s_{l})$. Combining this with the IPO loss (8), we recover Distilled DPO (Eq. 7 in Fisch et al. 2024). Using the same soft labels, but with the CE loss recovers RPO (see Sec. 3.3.2 in Adler et al. 2024). We provide more details on both algorithms in Appendix D. These natural extensions further demonstrate that our classification framework is fairly general as well as sufficiently flexible to capture a large number of existing DPO-style algorithms.
4 A Pitfall of DPO-style Algorithms
We now leverage the classification framework developed in Section 3 to explain a peculiar behavior of several DPO-style algorithms, namely the decrease, or even the collapse (to 00), of the winner response probability. We then use this insight to derive algorithms that address this phenomenon in Section 5. Several prior work (Adler et al., 2024; Xu et al., 2024; Pal et al., 2024; Fisch et al., 2024; Xiao et al., 2024) reported this undesirable behavior, empirically showed that it could result in suboptimal alignment, and proposed loss functions (algorithms) to address it.
Reinspecting the classification framework allows us to clearly identify the root-cause of this behavior as the underlying problem solved in these algorithms is under-specified. In practice, the under-specified nature of the problem often manifests itself in the form of collapsing the winner-loser probabilities to 00.
To understand this phenomenon, recall from Section 3 that the original problem can be reformulated as: $\min_{\theta}\sum_{\mathcal{D}}{\mathcal{L}}(p_{\theta},p)$. For simplicity, we use the soft labels $p=(1-\varepsilon,\varepsilon)$ for some $20<\varepsilon<1/$ 2. An optimal parameter $\theta$ is one that achieves 00 loss for all samples in $\mathcal{D}$, which means that $p_{\theta}(x,y_{w},y_{l})=p$, for all $(x,y_{w},y_{l})\in\mathcal{D}$. Setting $p_{\theta}^{w}(x,y_{w},y_{l})$ in (7) equal to $p^{w}=1-\varepsilon$, we obtain
$\pi_{\theta^{*}}(y_{w}|x)=\eta\cdot\pi_{\theta^{*}}(y_{l}|x)\ ,$ (9)
with $\eta\mathrel{\mathop{:}}=\sqrt[\beta]{(1-\varepsilon)/\varepsilon}\cdot\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}$. Note that the derivation using $y_{l}$ yields the same result. Thus, we have two probabilities, $\pi_{\theta^{*}}(y_{w}|x)$ and $\pi_{\theta^{*}}(y_{l}|x)$, that we aim to learn, but minimizing the loss only gives us one constraint specified in (9). This means that the original problem is under-specified. In Figure 1, we provide an illustration, and below we give a concrete example further highlighting that the two probabilities can move in arbitrary directions.
Figure 1: An illustration of the set of solutions that achieve 00 loss in DPO-style algorithms. The shaded gray is the set of feasible solutions. The red line passing the feasible set is defined by (9) and indicates the optimal set of solutions. All points on this line achieve 00 loss. Note that the case where the probabilities collapse to 0, i.e., $p_{\theta}(x,y_{w},y_{l})=(0^{+},0^{+})$ also lies on this line.
Remark 4.1 (A Concrete Example).
Suppose we have $\varepsilon=1/11$ and for simplicity we set $\beta=1$, $\pi_{\textrm{ref}}(y_{w}|x)=0.02$, and $\pi_{\textrm{ref}}(y_{l}|x)=0.01$, which result in $\eta=20$. We identify two pairs of probabilities that satisfy (9):
  1. 1.
[…]
  2. 2.
[…]
In the first case both probabilities (including the loser) increase under $\pi_{\theta^{*}}$ relative to $\pi_{\textrm{ref}}$. In sharp contrast, both probabilities (including the winner) decrease under $\pi_{\theta^{*}}$ relative to $\pi_{\textrm{ref}}$. Note that the increase is bounded as the two probabilities can go up until they hit $\pi_{\theta^{*}}(y_{w}|x)+\pi_{\theta^{*}}(y_{l}|x)=(1+\eta)\pi_{\theta^{*}}(y_{l}|x)=1$. Perhaps more concerning is the observation that the two probabilities can decrease arbitrarily and even collapse to 00 while still maintaining ${\mathcal{L}}\big{(}p_{\theta^{*}},p\big{)}=0$.
Note that for simplicity we worked with soft labels ($\varepsilon>0$). However, notice that with hard labels we often use early stopping, which in effect prevents the model from overfitting to the hard labels. Therefore, the problem remains under-specified even when working with hard labels.
5 Constrained Controlled Classification for DPO (C-3DPO)
We now present a general family of constraints to fix the under-specified nature of the DPO-style algorithms described in Section 4. These constraints are on the probability mass of the winner-loser pair and we use them to control how much this mass changes from the reference policy, $\pi_{\textrm{ref}}$, to target policy $\pi_{\theta}$. We then show how these constraints can be incorporated into any DPO-style loss function and propose our algorithms which we refer to as Constrained Controlled Classification for Direct Preference Optimization (C-3DPO).
The constraint, with respect to an arbitrary function $\varphi:\mathbb{R}\rightarrow\mathbb{R}$, takes the general form:
$\varphi\big{(}\pi_{\theta}(y_{w}|x)\big{)}$ $+\varphi\big{(}\pi_{\theta}(y_{l}|x)\big{)}$ (10)
$=\varphi\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}+\varphi\big{(}\pi_{\textrm{ref}}(y_{l}|x)\big{)}\ .$
Note that the right-hand-side is fixed during training, so the two probabilities on the left-hand-side cannot move in the same direction.
We now generalize this intuition by showing that when the constraint function $\varphi$ is monotonic and is added to the solution characterization of DPO-style algorithms (9), then we can control the direction of the movement of probability mass for winner-loser pairs.
Proposition 5.1.
Let $\varphi:\mathbb{R}\rightarrow\mathbb{R}$ be a monotonic function and assume that (10) holds. Then, $\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)$ and $\pi_{\theta^{*}}(y_{l}|x)<\pi_{\textrm{ref}}(y_{l}|x)$.
Proof.
For simplicity we assume that $\varphi$ is monotonically increasing (the same arguments can be easily applied to the monotonically decreasing case). From (9), we may write
$\frac{\pi_{\theta^{*}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}=\sqrt[\beta]{\frac{1-\varepsilon}{\varepsilon}}\cdot\frac{\pi_{\theta^{*}}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\ .$
Since $\varepsilon<1/2$, we have that $(1-\varepsilon)/\varepsilon>1$, and therefore
$\pi_{\theta^{*}}(y_{w}|x)>\frac{\pi_{\theta^{*}}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\cdot\pi_{\textrm{ref}}(y_{w}|x)\ .$ (11)
We will show that $\pi_{\theta^{*}}(y_{l}|x)<\pi_{\textrm{ref}}(y_{l}|x)$ by contradiction. To this end, we assume that $\pi_{\theta^{*}}(y_{l}|x)\geq\pi_{\textrm{ref}}(y_{l}|x)$. From (11), we immediately obtain that $\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)$. Applying the constraint $\varphi$ to the following two inequalities:
$\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)\quad\text{and}\quad\pi_{\theta^{*}}(y_{l}|x)\geq\pi_{\textrm{ref}}(y_{l}|x)\ ,$
and using its monotonicity, we obtain
$\varphi\big{(}\pi_{\theta^{*}}(y_{w}|x)\big{)}$ $>$ $\varphi\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}$
$\varphi\big{(}\pi_{\theta^{*}}(y_{l}|x)\big{)}$$\geq$ $\varphi\big{(}\pi_{\textrm{ref}}(y_{l}|x)\big{)}.$
By adding both sides of the above two inequalities, we get a contradiction to (10). Thus, proving that $\pi_{\theta^{*}}(y_{l}|x)<\pi_{\textrm{ref}}(y_{l}|x)$. Similarly, we can prove that $\pi_{\theta^{*}}(y_{w}|x)>\pi_{\textrm{ref}}(y_{w}|x)$, which concludes the proof. ∎
We just proved that any monotonic constraint function $\varphi$ guarantees that the learned policy, $\pi_{\theta^{*}}$, assigns a higher (lower) probability to the winner (loser) response than the one assigned to it by the reference policy $\pi_{\textrm{ref}}$. Thus, the resulting algorithm avoids the undesirable behavior of probability collapse for the winner response. It is now natural to ask what specific $\varphi$ we should use in the context of this constraint. We present two interesting candidates next.
5.1 Logarithmic Constraint $\;\varphi(x)\mathrel{\mathop{:}}=\log x$
Our first choice is to employ the logarithmic constraint:
$\log\left(\pi_{\theta}(y_{w}|x)\right)+\log\left(\pi_{\theta}(y_{l}|x)\right)=$ (12)
$\log\left(\pi_{\textrm{ref}}(y_{w}|x)\right)+\log\left(\pi_{\textrm{ref}}(y_{l}|x)\right)\ ,$
which is nice to work with empirically in light of the fact that all terms are in the log-space. Moreover, these log probabilities are already computed in DPO, which would help with efficient implementation of the corresponding C-3DPO algorithm.
Rather than using hard constraints, it is easier to compute the deviation from the constraint using either $\ell_{1}$ or $\ell_{2}$ norm, and then add it as a regularizer to the original DPO-style loss with a regularization parameter, $\lambda$, that trades-off the relative importance of the two terms. Equipping the DPO loss (5) with the logarithmic constraint (12), we obtain the following loss for C-3DPO:
$\min_{\theta}\;\sum_{\mathcal{D}}\;-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)$
$\;\;+\lambda\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}+\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)^{2}\ .$ (13)
In contrast to the hard constraint, in this case we do not necessarily force the winner (loser) probability to go up (down). Rather, we impose a penalty when the learner violates the constraint. Notice also that we added the penalty term to the original loss of DPO for simplicity, but in principle, the penalty term can be added to any DPO-style loss covered in our classification framework.
Further, we can show that employing the logarithmic constraint (12) has a meaningful RLHF interpretation. Recall that Rafailov et al. (2023) defined
$\hat{r}_{\theta}(x,y):=\beta\log\big{(}\pi_{\theta}(y|x)/\pi_{\textrm{ref}}(y|x)\big{)},\;\forall y\in\mathcal{Y},$ (14)
as implicit reward learned during DPO training. Using this notation, we can rewrite the objective (13) simply as
$-\log\sigma\big{(}\hat{r}_{\theta}(x,y_{w})-\hat{r}_{\theta}(x,y_{l})\big{)}+\frac{\lambda}{\beta^{2}}\big{(}\hat{r}_{\theta}(x,y_{w})+\hat{r}_{\theta}(x,y_{l})\big{)}^{2}\ .$
Under $\varphi(x)=\log x$, we solve the original RLHF problem akin to DPO, but we also incentivize that the sum of the implicit rewards for the winner and loser to stay around zero. It follows that the constraint regularizes the implicit rewards so as to avoid rewards that are (a) very large and (b) have the same sign. These two properties cannot co-exist when employing $\varphi(x)=\log x$, since doing so would yield a large magnitude inside the square and ultimately a large magnitude in the second piece of the loss. Intuitively this can hedge against probability collapse, because in the case of collapse both implicit rewards are large in magnitude and also both have negative sign, which the constraint will highly penalize.
5.2 Identity Constraint $\;\varphi(x)\mathrel{\mathop{:}}=x$
A second interesting choice would be to simply use the identity constraint:
$\pi_{\theta}(y_{w}|x)+\pi_{\theta}(y_{l}|x)=\pi_{\textrm{ref}}(y_{w}|x)+\pi_{\textrm{ref}}(y_{l}|x)\ .$ (15)
While (15) is also a plausible constraint, at first glance it is unclear how to implement it since the constraint is no longer in the log-space and is specified in terms of raw probabilities. Working with raw probabilities is prone to numerical underflow issues, thus, we would like to derive a constraint which is equivalent to (15) and operates in the log-space. To do so, we make use of the following lemma whose proof is reported in Appendix E.
Lemma 5.1.
For any two numbers $a$ and $b$, we have
$\log(a+b)=\log a-\log\sigma(\log a-\log b)\ .$
Applying $\log$ to both sides of (15), we obtain
$\log\big{(}\pi_{\theta}(y_{w}|x)+\pi_{\theta}(y_{l}|x)\big{)}=\log\big{(}\pi_{\textrm{ref}}(y_{w}|x)+\pi_{\textrm{ref}}(y_{l}|x)\big{)}\ ,$
which can be rewritten using Lemma 5.1 as
$\log\big{(}$ $\pi_{\theta}(y_{w}|x)\big{)}-\log\sigma\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)$ (16)
=\;= $\log\big{(}\pi_{\textrm{ref}}(y_{w}|x)\big{)}-\log\sigma\left(\log\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\ .$
Moving from (15) to (16), we have rewritten the constraint entirely in the log-space, thus, avoiding numerical issues, and similar to the logarithmic constraint in Section 5.1, allowing a straightforward implementation of the corresponding C-3DPO algorithm. Equipping the DPO loss (5) with the logarithmic constraint (16), we obtain the following loss for C-3DPO:
$\min_{\theta}\;\sum_{\mathcal{D}}\;-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)$
$\;\;+\lambda\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}+\log\sigma\left(\log\frac{\pi_{\textrm{ref}}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)\right.$
$\;\;\left.-\log\sigma\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\theta}(y_{l}|x)}\right)\right)^{2}\ .$ (17)
We conclude the section by noting that unlike $\varphi(x)=\log x$, deriving an RLHF interpretation under $\varphi(x)=x$ is subtle. That said, an interesting property under $\varphi(x)=x$ is that the winner probability can increase only by an amount equal to $\pi_{\textrm{ref}}(y_{l}|x)$. This means that we will not put the entire probability mass behind $y_{w}$. Thus, this constraint can serve as an additional guardrail to maintain a stochastic $\pi_{\theta}$.
6 Experiments
Figure 2: Head to head win-rate for four different implementations of C-3DPO against DPO. For each individual plot, we take the prompts from the held-out test set of Ultrafeedback Binarized and generate responses from the C-3DPO and DPO model. Then we present the prompt and two responses to Anthropic Claude Sonnet-3.5-v2 and ask which of the two answers is more helpful and honest. We then compute the win rate. Results are averaged over 10 different random seeds.
Figure 3: A comparison between C-3DPO and baselines. (Left) The two variants of C-3DPO-Log better align Zephyr-7b-SFT relative to vanilla DPO. (Right) C-3DPO variants also improve upon existing baselines that are designed to mitigate the collapse issue of DPO.
In this section, we empirically evaluate C-3DPO across 2 standard datasets and 3 initial checkpoints.
6.1 Ultrafeedback Binarized
The original Ultrafeedback dataset comprises 64k prompts, each paired with four completions from various models. To construct Ultrafeedback-Binarized, GPT-4 assigns scores based on criteria such as helpfulness. UltraFeedback_binarized was then derived by selecting the highest-scoring completion as $y_{w}$ and a randomly selected as $y_{l}$.
Figure 4: A comparison between C-3DPO and baselines with the GPT-J model on the TL;DR dataset.
Following Rasul et al. (2024), we align Zephyr-7B-SFT by performing various preference optimization algorithms on this initial checkpoint. To holistically evaluate the final model,  Rasul et al. (2024) used MT-Bench, a multi-turn benchmark that uses GPT-4 to judge models’ performance in $8$ different categories: Writing, Roleplay, Reasoning, Math, Coding, Extraction, STEM, and Humanities. In addition to MT-Bench, we also present the head to head win rate of C-3DPO against DPO by using 200 prompts from the held-out test set of Ultrafeedback-Binarized. We compute the win-rate by asking Anthropic’s Claude Sonnet-3.5-v2 which response is more helpful. The exact prompt used for Claude is in Appendix G.
Figure 5: Comparison between C-3DPO and DPO for aligning the Olmo-13B-SFT model using the Ultrafeedback-Binarized dataset.
Figure 6: Comparing C-3DPO and the baseline on the TL;DR dataset. The first four plots show the win rates of checkpoints across the first epoch against the preferred human summaries. The last figure shows the win rates of all checkpoints at different temperatures.
Recall from Section 5 that we proposed two candidates for the constraint function $\varphi$, namely the logarithmic and identity functions. Moreover, notice from (17) that we measure the deviation from the constraint using an $\ell_{2}$ penalty. Alternatively, we can measure this deviation using an $\ell_{1}$ penalty. Therefore, together, we have four specific implementations of C-3DPO, namely: C-3DPO-Log-$\ell_{1}$, C-3DPO-Log-$\ell_{2}$, $-\mathcal{I}$-$\ell_{1}$, and $-\mathcal{I}$-$\ell_{2}$. In Appendix F, we present a pseudo-code for each of these 4 variations. In Figure 2, we show the head to head win rate of each of these 4 implementations against DPO. For all 4 implementations we used the hyper-parameter $\lambda=2\times 10^{-4}$. While we did not tune $\lambda$ for each of the 4 implementations separately, doing so may increase the gap between C-3DPO and DPO.
From this result, it is clear that C-3DPO-Log is resulting in better alignment. Therefore, we use the two final checkpoints from C-3DPO-Log for MT-Bench evaluation. In Figure 3 we show the MT-Bench evaluation for C-3DPO against vanilla DPO, as well as related DPO-style algorithms that discuss the collapse of probabilities in DPO and aim to mitigate it, such as Cal-DPO (Xiao et al., 2024), SPPO (Wu et al., 2024), and DPOP (Pal et al., 2024). C-3DPO is the most competitive algorithm.
We then use C-3DPO-Log on a larger initial checkpoint, namely the Olmo-13B-SFT model from AllenAI (OLMo et al., 2024). We use the same Ultrafeedback-Binarized dataset to align this initial SFT model, and compare C-3DPO-Log against DPO. In Figure 5, we again see the superiority of C-3DPO-Log against vanilla DPO suggesting that these improvements may scale to larger models.
6.2 Reddit TL;DR
We then evaluate our proposed method on a summarization task with human-assigned scores for pairs of summaries. For this purpose, we employ the Reddit TL;DR dataset from Stiennon et al. (2022), specifically focusing on the posts under relationships and relationship_advice subreddits. We follow Amini et al. (2024) in creating the dataset and include pairs of summaries where one received a higher quality score than the other. The training and validation dataset has 572 and 102 rows respectively, each containing one Reddit post and four to twenty-four user generated summaries and their corresponding scores. During training and testing, we select the highest-scoring summary as $y_{w}$ and a randomly selected alternative as $y_{l}$.
We align the GPT-J developed by Wang and Komatsuzaki (2021) with vanilla DPO as well as C-3DPO algorithms. Specifically, we first run one SFT epoch with the Reddit TL;DR dataset on the GPT-J checkpoint, where we maximize the probability of generating the winner summary, and then take the SFT checkpoint for subsequent alignment experiments. We evaluate the final checkpoints by computing win rates against $y_{w}$ following Rafailov et al. (2023); Amini et al. (2024). We use Claude Sonnet 3.5 v2 as a judge to determine which summary is more concise and precise, and provide prompt used for Claude in the Appendix H. As shown in Figure 6, our proposed method improves upon the baseline when using higher temperatures.
7 Related Work
A few recent papers have presented a unifying perspective on DPO-style algorithms. Notably, Tang et al. (2024) presented a generalization of DPO where different supervised learning losses are applied to the difference of implicit rewards, $\big{(}\hat{r}(x,y_{w})-\hat{r}(x,y_{l})\big{)}/\beta$. In contrast, we make a clear connection between DPO-style algorithms and classification, and also extend our results to lists and auxiliary information. A second notable example was to generalize from KL to any f-divergence when measuring the discrepancy between the target and reference model (Han et al., 2024). We also note that the first ingredient of our classification framework, namely $\pi_{\theta}$, was used by Sharifnassab et al. (2024) to propose a soft version of DPO.
As mentioned in Section 4, earlier works studied the undesirable decrease of the winner probability during training. These works explained this phenomenon from different angles, and thus, proposed different losses to address it. Here we provide a brief overview of a number of these results, especially those that we experimentally compare against. Pal et al. (2024) proposed DPOP which addresses the phenomenon by adding the penalty term $\max\big{(}0,-\hat{r}(x,y_{w})/\beta\big{)}$ within the log-sigmoid of the DPO loss (5). DPOP can also be viewed as DPO with a modified BT model. In this sense, it has similarities with the $\alpha$-DPO loss (Shen et al., 2024).
Xiao et al. (2024) attributed the undesirable behavior to the contrastive loss of DPO not being scale-calibrated, i.e., ignoring the absolute values of implicit rewards, $\hat{r}(x,y_{w})$ and $\hat{r}(x,y_{l})$. They address this by constraining the implicit rewards to a scale that matches the ground-truth rewards $r(x,y)$. Thus, in their proposed loss, Cal-DPO, they add the square-loss $\big{(}\hat{r}(x,y)-r(x,y)\big{)}^{2},\;y\in\{y_{w},y_{l}\}$ to the DPO loss (without $\beta$). Of course, since they do not have access to the ground-truth reward, they replace it with $1/2$ and $-1/2$ for $y_{w}$ and $y_{l}$, respectively. A loss very similar to Cal-DPO was proposed by Wu et al. (2024) and they named it SPPO. It is simply Cal-DPO without the DPO loss. Finally, D’Oosterlinck et al. (2024) proposed APO, which offer fine-grained control over the implicit rewards.
8 Conclusion & Future Work
In this work, we made two contributions to the preference optimization literature. First, we showed that many DPO-style algorithms can be viewed as classification algorithms. In this classification framework, one may choose a specific classification label (soft or hard, using auxiliary information or not) and a specific loss to obtain a DPO-style algorithm. We believe that revealing this connection promotes a mutual transfer of ideas between the classification and preference optimization.
The classification framework also revealed that the underlying problem solved in DPO-style algorithms is under-specified and therefore susceptible to the collapse of both winner-loser probabilities. By endowing the DPO-style algorithms with our newly proposed constraints, we showed that the problem becomes specified, and thus, probability-collapse can be avoided.
A curious observation we made is that while we do want to avoid probability collapse, it is also not true that increasing the probability of the winner will always lead to the best final model. This means that the best performance is usually achieved by increasing the probability of some of the unseen answers. Thus, in future, it is interesting to do a more systematic study of the kinds of unseen responses for which DPO-style algorithms increase the probability during training.
Appendix A Derivations of IPO as a Classification Algorithm
We recall that IPO can be formulated as a classification problem with the soft labels $p\mathrel{\mathop{:}}=(p^{w},p^{l})=(\sigma(1/2),\sigma(-1/2))$ and the loss given by
$\mathcal{L}\big{(}p_{\theta},\ p\big{)}=\left(\log\frac{p^{w}_{\theta}}{p^{l}_{\theta}}-\log\frac{p^{w}}{p^{l}}\right)^{2}\ .$ (18)
To show that we indeed recover the IPO, we first note that
$\frac{p^{w}}{p^{l}}=\frac{\sigma(1/2)}{\sigma(-1/2)}=\frac{1+\exp(1/2)}{1+\exp(-1/2)}=\exp(1/2)\ ,$
where the second equality follows from the definition of the sigmoid $\sigma(x)=1/(1+\exp(-x))$. Moreover, using the definitions of $p^{w}_{\theta}$ (see (7)) and $p^{l}_{\theta}$, we obtain that
$\frac{p^{w}_{\theta}}{p^{l}_{\theta}}=\frac{\big{(}\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\textrm{ref}}(y_{w}\mid x)}\big{)}^{\beta}}{\big{(}\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\textrm{ref}}(y_{l}\mid x)}\big{)}^{\beta}}=\left(\frac{\pi_{\theta}(y_{w}|x)\pi_{\textrm{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)\pi_{\textrm{ref}}(y_{w}|x)}\right)^{\beta}\ .$ (19)
Plugging these two developments to the loss in (18) yields
$\mathcal{L}\big{(}p_{\theta},\ p\big{)}=\left(\log\frac{p^{w}_{\theta}}{p^{l}_{\theta}}-\log\frac{p^{w}}{p^{l}}\right)^{2}=\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\textrm{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)\pi_{\textrm{ref}}(y_{w}|x)}-\frac{1}{2}\right)^{2}=\left(\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\textrm{ref}}(y_{l}|x)}{\pi_{\theta}(y_{l}|x)\pi_{\textrm{ref}}(y_{w}|x)}-\frac{1}{2\beta}\right)^{2}\ ,$
which is exactly IPO (see Eq. (17) of Azar et al. 2024).
Appendix B CDPO
Recall the CDPO [Mitchell, 2024] loss is given for any triplet $(x,y_{w},y_{l})$ by
$-(1-\varepsilon)\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}\right)-\varepsilon\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}-\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}\right)\ ,$
and then summed over all the triplets in the dataset ${\cal D}$. In order to see CDPO as a classification with the soft labels $p:=(p^{w},p^{l})=(1-\varepsilon,\varepsilon)$ and the CE loss, we will use the following technical fact
$\sigma(\beta\log(a/b))=\frac{1}{1+\exp(-\beta\log(a/b))}=\frac{1}{1+\frac{\exp{\log b^{\beta}}}{\exp{\log a^{\beta}}}}=\frac{1}{1+\frac{b^{\beta}}{a^{\beta}}}=\frac{a^{\beta}}{a^{\beta}+b^{\beta}}.$
Therefore, with $a=\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}$ and $b=\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}$ we get from (7) that $\sigma(\beta\log a/b)=p^{w}_{\theta}$. Similarly, we get that $\sigma(\beta\log(b/a))=p^{l}_{\theta}$. Therefore, the CDPO loss can be written as follows
$-(1-\varepsilon)\log p^{w}_{\theta}-\varepsilon\log p^{l}_{\theta}=-p^{w}\log p^{w}_{\theta}-p^{l}\log p^{l}_{\theta}\ ,$
which is exactly the CE loss on the vectors $p_{\theta}=(p^{w}_{\theta},p^{l}_{\theta})$ and $p=(p^{w},p^{l})$.
Appendix C DPO (PL)
In this setting, we are given a dataset of the form $\mathcal{D}=\{(x,y_{1},y_{2},\ldots,y_{N})\}$. In order to recover, the DPO with Plackett-Luce [Rafailov et al., 2023], we need to generalize the definition of the probability vector $p_{\theta}$ from pairs as in (7) to the following $N-1$ subsets of the list, namely the first $N$, then the first $N-1$, then first $N-2$ and so on. More precisely, for any $1\leq n<N$ we define
$p_{\theta}(x,y_{n},y_{n+1},\ldots,y_{N})=\textrm{softmax}\big{(}\left(r_{\theta}(x,y_{n}),r_{\theta}(x,y_{n+1}),\ldots,r_{\theta}(x,y_{N})\right)\big{)}\ .$
In this case, the hard label vectors are defined, for any $1\leq n<N$, by $p^{[n,N]}:=(1,0,\ldots,0)\in\mathbb{R}^{N-n+1}$. Now, using the CE loss we get the desired result as follows
$-\sum_{n=1}^{N-1}\sum_{i=n}^{N}p_{i}^{[n,N]}\log p_{\theta}^{i}(x,y_{n},y_{n+1},\ldots,y_{N})$ $=$ $-\sum_{n=1}^{N-1}\log\frac{\exp\left(\beta\log\frac{Z(x)\pi_{\theta}(y_{n}|x)}{\pi_{\textrm{ref}}(y_{n}|x)}\right)}{\sum_{i=n}^{N}\exp\left(\beta\log\frac{Z(x)\pi_{\theta}(y_{i}|x)}{\pi_{\textrm{ref}}(y_{i}|x)}\right)}$
$=$ $-\sum_{n=1}^{N-1}\log\frac{\exp\left(\beta\log\frac{\pi_{\theta}(y_{n}|x)}{\pi_{\textrm{ref}}(y_{n}|x)}\right)}{\sum_{i=n}^{N}\exp\left(\beta\log\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\textrm{ref}}(y_{i}|x)}\right)}$
$=$ $-\log\prod_{n=1}^{N-1}\frac{\exp\left(\beta\log\frac{\pi_{\theta}(y_{n}|x)}{\pi_{\textrm{ref}}(y_{n}|x)}\right)}{\sum_{i=n}^{N}\exp\left(\beta\log\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\textrm{ref}}(y_{i}|x)}\right)}\ .$
Appendix D RPO and Distilled DPO
As we discussed in Section 3.2, RPO can be reformulated as a classification with soft labels, which are defined by $p=\text{softmax}(s_{w},s_{l})$. Therefore, we immediately see that
$p^{w}=\frac{\exp(s_{w})}{\exp(s_{w})+\exp(s_{l})}=\frac{1}{1+\exp(-(s_{w}-s_{l}))}\ ,$
and thus $p^{w}=\sigma(s_{w}-s_{l})$. Similarly, we get that $p^{l}=\sigma(-(s_{w}-s_{l}))$. Thus, RPO can be seen as a generalization of CDPO where the soft labels are given by a certain score and not fixed. To recover the RPO loss we denote $a=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\textrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\textrm{ref}}(y_{l}|x)}$ and $b=s_{w}-s_{l}$. Then, we get that
$\sigma(b)\log\frac{\sigma(b)}{\sigma(a)}+(1-\sigma(b))\log\frac{1-\sigma(b)}{1-\sigma(a)}=p^{w}\log\frac{p^{w}}{p_{\theta}^{w}}+p^{l}\log\frac{p^{l}}{p_{\theta}^{l}}\ .$
By eliminating the constant terms (with respect to $\theta$) $p^{w}\log p^{w}+p^{l}\log p^{l}$, we indeed get the CE loss.
To recover the Distilled DPO (Equation (7) of  Fisch et al. [2024]), we consider the soft labels $p=\text{softmax}(s_{w},s_{l})$ with the loss of (8)
$\mathcal{L}\big{(}p_{\theta},\ p\big{)}=\left(\log\frac{p^{w}_{\theta}}{p^{l}_{\theta}}-\log\frac{p^{w}}{p^{l}}\right)^{2}\ .$
Indeed, it this case we have that
$\frac{p^{w}}{p^{l}}=\frac{\exp(s_{w})}{\exp(s_{w})+\exp(s_{l})}\cdot\frac{\exp(s_{w})+\exp(s_{l})}{\exp(s_{l})}=\frac{\exp(s_{w})}{\exp(s_{l})}\ ,$
and therefore $\log(p^{w}/p^{l})=s_{w}-s_{l}$. Combining this with (19) yields the desired result.
Appendix E Proof of Lemma 5.1
First, we recall the lemma. See 5.1
Proof.
First, we write
$a+b=a\left(\frac{a}{a+b}\right)^{-1}\ ,$
which tanks to classical logarithmic rules yields that
$\log(a+b)=\log a-\log\frac{a}{a+b}=\log a-\log\frac{1}{1+b/a}\ .$
Using the definition of the sigmoid function we get
$\frac{1}{1+b/a}=\frac{1}{1+\exp(\log b-\log a)}=\sigma(\log a-\log b)\ ,$
which proves the desired result. ∎
Appendix F C-3DPO implementation details
We show implementation details of C-3DPO-Log-$\ell_{1}$, C-3DPO-Log-$\ell_{2}$, $-\mathcal{I}$-$\ell_{1}$, and $-\mathcal{I}$-$\ell_{2}$ below.
def dpo_loss(pi_logps, ref_logps, yw_idxs, yl_idxs, beta, algo_name, reg_coeff):
    pi_logps: policy logprobs, shape (B,)
    ref_logps: reference model logprobs, shape (B,)
    yw_idxs: preferred completion indices in [0, B-1], shape (T,)
    yl_idxs: dispreferred completion indices in [0, B-1], shape (T,)
    beta: temperature controlling strength of KL penalty
    Each pair of (yw_idxs[i], yl_idxs[i]) represents the indices of a single
    preference pair.
    pi_yw_logps,  pi_yl_logps =  pi_logps[yw_idxs],  pi_logps[yl_idxs]
    ref_yw_logps, ref_yl_logps = ref_logps[yw_idxs], ref_logps[yl_idxs]
    pi_logratios  = pi_yw_logps - pi_yl_logps
    ref_logratios = ref_yw_logps - ref_yl_logps
    losses = -F.logsigmoid(beta * (pi_logratios - ref_logratios))
    if algo_name == ’c3dpo_log_l1’:
        reguralized_losses_without_square = \
            pi_yw_logps + pi_yl_logps - ref_yw_logps - ref_yl_logps
        reguralized_losses = (reguralized_losses_without_square) ** 2
        losses = losses + reg_coeff * reguralized_losses
    elif algo_name == ’c3dpo_log_l2’:
        reguralized_losses = F.l1_loss(
            pi_yw_logps + pi_yl_logps,
            ref_yw_logps + ref_yl_logps
        losses = losses + reg_coeff * reguralized_losses
    elif algo_name == ’c3dpo_i_l1’:
        reguralized_losses_without_square = \
            pi_yw_logps - F.logsigmoid(pi_yw_logps - pi_yl_logps) - \
                (ref_yw_logps - F.logsigmoid(ref_yw_logps - ref_yl_logps))
        reguralized_losses = (reguralized_losses_without_square) ** 2
        losses = losses + reg_coeff * reguralized_losses
    elif algo_name == ’c3dpo_i_l2’:
        reguralized_losses = F.l1_loss(
            pi_yw_logps - F.logsigmoid(pi_yw_logps - pi_yl_logps),
            ref_yw_logps - F.logsigmoid(ref_yw_logps - ref_yl_logps)
        losses = losses + reg_coeff * reguralized_losses
    rewards = beta * (pi_logps - ref_logps).detach()
    return losses, rewards
Appendix G Ultrafeedback Binarized Claude 3.5 Sonnet v2 win rate prompt and hyperparameters
In this section we include the prompt used to generate win rates for the Ultrafeedback Binarized experiments. We use Claude Sonnet 3.5 v2 (AWS Bedrock model ID anthropic.claude-3-5-sonnet-20241022-v2:0) to generate win rates. We set the max_tokens to 1024, temperature to 0, and used default value 0.999 for top_p and top_k disabled.
For the following query to a chatbot, which response is more helpful?
Query: <prompt>
Response A:
<one of the responses>
Response B:
<the other response>
FIRST provide a one-sentence comparison of the two responses and explain
which you feel is more helpful. SECOND, on a new line, state only "A" or "B"
to indicate which response is more helpful. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
More helpful: <"A" or "B">
Appendix H Reddit TL;DR Claude 3.5 Sonnet v2 win rate prompt and hyperparameters
In this section we include the prompt used to generate win rates for the Reddit TL;DR experiments. We use Claude Sonnet 3.5 v2 (AWS Bedrock model ID anthropic.claude-3-5-sonnet-20241022-v2:0) to generate win rates. We set the max_tokens to 1024, temperature to 0, and used default value 0.999 for top_p and top_k disabled.
Which of the following summaries does a better job of summarizing the most
important points in the given forum post, without including unimportant or
irrelevant details? A good summary is both precise and concise.
Post: {prompt}
Summary A:
{baseline_response}
Summary B:
{generated_response}
FIRST provide a one-sentence comparison of the two summaries, explaining
which you prefer and why. SECOND, on a new line, state only "A" or "B" to
indicate your choice. Your response should use the format:
Comparison: <one-sentence comparison and explanation>
Preferred: <"A" or "B">
Appendix I Additional win rates analysis of C-3DPO with Zephyr-7B-SFT aligned on Ultrafeedback Binarized
Figure 7: Win rates comparison of Zephyr-7B-SFT aligned on Ultrafeedback Binarized using DPO and C-3DPO. The first 4 plots show win rates of C-3DPO at individual checkpoints across 10 different inference runs. The last plot shows mean and standard error of win rates across all checkpoints and all inference runs.
Figure 7 shows win rates comparison between Zephyr-7B-SFT aligned with DPO and C-3DPO. We align Zephyr-7B-SFT following Rasul et al. [2024] using DPO, C-3DPO-Log-$\ell_{1}$, C-3DPO-Log-$\ell_{2}$, $-\mathcal{I}$-$\ell_{1}$, and $-\mathcal{I}$-$\ell_{2}$ for one epoch, all C-3DPO use hyper-parameter $\lambda=2\times 10^{-4}$. With each checkpoint, we generate responses using test split of Ultrafeedback Binarized using hyperparameters max_tokens=1000, temperature=1.0, top_p=0.9, top_k=50. Different from the head to head setting, we ask Claude to compare the generated response directly with the preferred response in the dataset. The win rates and standard errors are calculated based on 10 different inference runs.